# RIVCoin: an alternative, integrated, CeFi/DeFi-Vaulted Cryptocurrency

### v1

@@ 9 Conclusion
In this Whitepaper, we introduced RIVCoin, a cryptocurrency stabilized by a combination of CeFi and DeFi assets, fully compliant with all the relevant laws and regulations worldwide.
RIVCoin will be accessible to all Users via subsequent Token Mintings. By buying RIVCoin and providing liquidity, Users can earn rewards on their investment, allowing them to build a second income stream, and being able at the same time to spend it and use it. We, as RIV Capital, believe that Blockchain and DeFi will democratize finance allowing a fairer redistribution of wealth, co-operating and integrating with the traditional universe.
@@
We have explored the problem from a different point of view. Let’s bear in mind this analogy: as well as we have to be creators and not competitors, we found the way to create and add wealth instead of causing its dispersion. No doubt that such a cryptocurrency will increase the capacity of the wealthy people, but the marginal utility is bigger for Retail Users who will benefit, with the placet of Wealthier Users, from the bigger amounts invested by these ones. The last ones look for stability, the first ones would like not to be frauded participating at a fair game.
The Blockchain Revolution will enable a fairer world [18] where everyone can profit. We believe that anyone has the right to live a life at its fullest potential: by creating RIVCoin, we want to allow you to achieve this objective.

### v2

@@ 6 Conclusion
We have explored the problem from a different point of view. Let’s bear in mind this analogy: as well as we have to be creators and not competitors, we found the way to create and add wealth instead of causing its dispersion.
No doubt that such a cryptocurrency will increase the capacity of the wealthy people, but the marginal utility is bigger for Retail Users who will benefit, with the placet of Wealthier Users, from the bigger amounts invested by these ones. The last ones look for stability, the first ones would like not to be frauded participating at a fair game.
The Blockchain Revolution will enable a fairer world [bcstatistics] where everyone can profit. We believe that anyone has the right to live a life at its fullest potential: by creating RIVCoin, we want to allow you to achieve this objective by, to some extend simply “leader following”.
[…]